OpenStudy (anonymous):

Can someone give me an example of arational equation that has an extraneous solution using the following formula?

OpenStudy (anonymous):

\[\frac{ x+a }{ ax } = \frac{ b }{ x }\]

OpenStudy (anonymous):

@GeniousCreation

OpenStudy (anonymous):

1s this making any sense to you?

OpenStudy (anonymous):

Because its not making sense to me.

OpenStudy (anonymous):

It's a little difficult. But is making a little sense. So yeah...

OpenStudy (anonymous):

1 know what an extraneous solution is 1 just don't know how to get one...

OpenStudy (anonymous):

Does b have to be the extraneous solution?

OpenStudy (anonymous):

dont think so

OpenStudy (anonymous):

k.

OpenStudy (anonymous):

1 have to go eat dinner but 1 should be back in a bit. Let me know if you figure anything out.

OpenStudy (anonymous):

Perhaps, the variables have to match so: x=5 a=1 b=6 \[\frac{ 5+1 }{ 5(1) }=\frac{ 6 }{ 5 }\] l these are dependent on one another. Hope this helps!

OpenStudy (anonymous):

@bloofoffiction : I'm not entirely sure. Wait for a few mins, but if it is really needed, you can write it down.

OpenStudy (anonymous):

back

OpenStudy (anonymous):

1 think the x in the equation is a variable.

OpenStudy (anonymous):

yeah it is.

OpenStudy (anonymous):

how would you solve it?

OpenStudy (anonymous):

I think I already solved it.

OpenStudy (asnaseer):

Your equation was:\[\frac{ x+a }{ ax } = \frac{ b }{ x }\tag{1}\]If you cross-multiply the terms you get:\[x(x+a)=abx\]Which leads to:\[x(x+a)-abx=0\]\[\therefore x(x+a-ab)=0\]\[\therefore x=0\text{ or }x+a-ab=0\]But \(x=0\) is not valid as that would imply that equation (1) would be dividing by zero. Therefore the only valid solution is:\[x+a-ab=0\]\[\therefore x=ab-a=a(b-1)\]

OpenStudy (anonymous):

0.o

OpenStudy (asnaseer):

I am assuming you were trying to solve for \(x\)?

OpenStudy (anonymous):

1'm trying to find the extraneous solution

OpenStudy (asnaseer):

The extraneous solution here was \(x=0\)

OpenStudy (anonymous):

"Part 1. Show all work to solve for x in the equation and check the solution. Part 2. Explain how to identify the extraneous solution and what it means."

OpenStudy (asnaseer):

It is "extraneous" because it would be mean equation (1) involves dividing by zero which is not a valid operation

OpenStudy (anonymous):

okay so the extraneous solution is zero and x is 5?

OpenStudy (asnaseer):

where did you get x=5 from?

OpenStudy (anonymous):

@GeniousCreation

OpenStudy (asnaseer):

Did you follow my reasoning above where I deduced that:\[x=a(b-1)\]

OpenStudy (anonymous):

You were not here when 1 was given that answer.

OpenStudy (asnaseer):

Sorry - I don't understand what you are trying to say?

OpenStudy (anonymous):

Okay scratch that. Can you give me an equation following the specifications you've made?

OpenStudy (asnaseer):

It might help if you posted the whole of the question you were given here

OpenStudy (anonymous):

Task 4—Solving Rational Equations Using the equation below as a model, fill in numbers in the place of a and b to create a rational equation that has an extraneous solution. = Part 1. Show all work to solve for x in the equation and check the solution. Part 2. Explain how to identify the extraneous solution and what it means.

OpenStudy (anonymous):

\[\frac{ x+a }{ ax } = \frac{ b }{ x }\]

OpenStudy (asnaseer):

Ok, so I showed you how this leads to:\[x=a(b-1)\]and how we removed the extraneous solution \(x=0\) and the reasons for this. We can therefore pick any values for \(a\) and \(b\) as long as it does not lead to \(x=0\). So, for example, we cannot pick \(b=1\) as this leads to \(x=0\). So, lets pick \(a=10\) and \(b=5\) for example to get:\[x=a(b-1)=10(5-1)=40\]

OpenStudy (anonymous):

okay

OpenStudy (asnaseer):

Putting this back into our original equation leads to:\[\frac{x+a}{ax}=\frac{b}{x}\]\[\therefore \frac{40+10}{10\times40}=\frac{5}{40}\]\[\therefore \frac{50}{400}=\frac{5}{40}\]which is valid so our solution here works.

OpenStudy (anonymous):

awesome.

OpenStudy (asnaseer):

Please let me know if you need any further explanations - I would be more than happy to help :)

OpenStudy (anonymous):

You are currently my favorite person. Arigato!

OpenStudy (asnaseer):

yw my friend! :)